# PRISM: Proof-Carrying Artifact Generation through LLM $\times$ MDE Synergy and Stratified Constraints

### v1

@@ 4.5. RQ3: Cross-Domain Transferability via UMM S2 Path
We evaluate the framework in a domain without an explicit meta-model by exercising the UMM S2 path (inductive meta-model construction from natural-language sources). This cross-domain probe uses Private International Law (PIL) jurisdiction determination under the Brussels I bis Regulation as a representative case study, and asks whether the UMM–ICM–CVG architecture can be transferred into a high-stakes legal reasoning setting that does not provide a pre-existing, formally specified metamodel. This experiment is intended as an architectural validation of applicability, not as an attempt to build a production-ready legal AI system or to claim comprehensive doctrinal coverage.
@@ 4.5.1. Experimental Setup
We examine a legal domain without a formal schema (Brussels I bis). Prior work notes domain hallucination risks (Magesh et al., 2025); we therefore use the UMM S2 path to induce entities/relations/constraints from text and test whether layered enforcement reduces doctrinal errors (Magesh et al., 2025).
@@
Figure 9. UMM S2 pipeline architecture for cross-domain Private International Law evaluation. The pipeline instantiates a UMM mini-instance from fact patterns, links induced factors to Brussels I bis provisions through ICM construction, applies promote-aware constrained decoding with JSON Schema enforcement, and performs multi-layer verification with deterministic repair. Output comprises Evidence Registry and Structured Decision JSON forming a complete Decision Package.:
@@
  • RAG: Retrieval-augmented prompting in the style of commercial legal assistants studied in (Magesh et al., 2025). The model receives the fact pattern plus top-$k$ Brussels I bis provisions retrieved for that scenario, but generation remains unconstrained: it may assert a jurisdictional basis even if statutory preconditions are not actually satisfied.
  • PRISM: Full UMM $\rightarrow$ ICM $\rightarrow$ CVG S2 pipeline. The system first induces the structured mini-instance (UMM), links factors to provisions (ICM), then performs promote-aware JSON Schema-constrained decoding followed by deterministic verification and repair. This configuration enforces precedence, exclusivity, pairing requirements, and formality gates before finalizing the decision.
@@
The evaluation uses 30 expert-authored Brussels I bis jurisdiction cases spanning the full hierarchy of bases of jurisdiction: exclusive (Article 24), agreement-based (Articles 25 and 31), lis pendens / first-seised stay (Article 31), special jurisdiction (Article 7), general jurisdiction (Article 4), and appearance (Article 26). Each case was drafted and doctrinally validated by a subject-matter expert in European civil procedure. As noted above, this dataset is intended to probe architectural transfer, not to claim full coverage of real-world PIL litigation.
@@ 4.5.2. Results
Table 7 summarizes results across the three configurations. PRISM improves correctness, citation faithfulness, structural compliance, and priority alignment relative to both Baseline and RAG. RAG improves over the Baseline in several dimensions but still fails to guarantee constraint compliance.
Table 7. Cross-domain evaluation on 30 Brussels I bis jurisdiction cases comparing Baseline (LLM-only), RAG retrieval augmentation, and the full PRISM pipeline.
Metric Baseline RAG PRISM
@@
PRISM achieves perfect Schema-Pass (1.000) and Rule-OK (1.000), meaning every final output is syntactically valid and satisfies precedence, exclusivity, and pairing constraints. Promotion Accuracy (alignment between the claimed forum type and the strongest legally available basis) increases from 0.367 (Baseline) to 0.567 (RAG) and 0.933 (PRISM), indicating that PRISM’s promote-aware decoding and deterministic repair suppress “everything applies” answers and force adherence to the Brussels I bis priority ladder.
Abstention Quality—the ability to defer judgment when required information is missing—is 0.500 for Baseline, 0.800 for RAG, and 0.800 for PRISM. RAG benefits here because retrieved provisions explicitly enumerate preconditions; the model can recognize that those preconditions are not satisfied and refuse to commit. PRISM preserves that epistemic caution while additionally enforcing that abstention is required when, for example, Article 25 formality prerequisites are underspecified.
@@
PRISM incurs higher cost due to multi-stage generation, schema-constrained decoding, and post-hoc verification. This overhead is the direct cost of guaranteeing structured legality.
@@ 4.5.3. Analysis and Discussion
These results show that the UMM–ICM–CVG architecture transfers to a non-engineering, text-governed domain (PIL jurisdiction) via the S2 inductive pathway. Without any pre-existing metamodel, the system can induce a structured mini-instance, align facts to candidate provisions, enforce hierarchical precedence, and output auditable decisions. PRISM delivers higher Legal-Correct@1 than both Baseline and RAG (0.467 vs. 0.200 / 0.300), far higher Promotion Accuracy, and perfect Schema-Pass / Rule-OK. This indicates that layered constraint enforcement meaningfully outperforms unconstrained prompting and improves upon retrieval augmentation alone (Magesh et al., 2025).
@@ Discussion.
RAG improves correctness (0.300 vs. 0.200 Baseline) and abstention (0.800 vs. 0.500), echoing (Magesh et al., 2025), but does not enforce doctrinal validity. PRISM’s UMM $\rightarrow$ ICM $\rightarrow$ CVG pipeline adds precedence rules, exclusivity constraints, and formality gates, yielding 0.933 Promotion Accuracy, 0.778 Citation Precision, and perfect structural legality (Schema-Pass = Rule-OK = 1.000). Each case logs UMM instance, factor-provision mappings, decoding trace, and repair actions, providing end-to-end provenance for practitioner review.
This 30-case Brussels I bis evaluation is a feasibility probe testing architectural transfer, not claiming exhaustive legal coverage or deployment readiness. Scaling requires broader doctrinal coverage, multi-jurisdictional support, and human oversight in professional workflows.
@@ 5. Conclusion and Future Work
This paper presented PRISM, a high-assurance generation framework integrating LLMs with model-driven engineering and formal methods to produce verifiable, auditable artifacts in safety- and regulation-critical domains. PRISM combines a Unified Meta-Model (UMM) for typed domain semantics, an Integrated Constraint Model (ICM) for constraint aggregation, and Constraint-guided Verifiable Generation (CVG) with Audit-Guided Repair (AGR) to produce verifiable artifacts $\mathfrak{A}=(a,\Pi,\varphi)$. Constraints are stratified across two layers: L1 enforces structural rules during decoding (Theorem 12), while L2 validates semantic/logic constraints post-generation with machine-checkable certificates.
@@
  • System-level semantics at scale. In RQ2, architectural drift and dangling cross-file references appear in more complex systems. PRISM can surface these violations and drive AGR, but does not yet guarantee global semantic invariants for arbitrarily dense dependency graphs.
  • Cost and latency. DFA/PDA masking with per-step audit logging introduces decoding overhead, and SHACL/SMT validation plus AGR adds post-generation latency. In regulated workflows this is acceptable, but embedded/real-time deployment (e.g., ECU configuration on-device) may require lighter-weight variants.
  • Human review remains essential. PRISM narrows the surface that humans must review and provides $\Pi$ and $\tau$ to justify remaining issues, but it does not claim unsupervised certification. Certified deployment still requires domain experts and regulatory sign-off.
@@
We outline several directions for advancing PRISM and the broader UMM–ICM–CVG paradigm:
@@
    Incremental semantic and logic validation. Beyond the post-artifact L2 validation described in Limitations, we aim to integrate streaming SHACL engines (Ye et al., 2023) and incremental SMT solvers accepting partial formulas, enabling Layer-2 feedback during generation. A key challenge is exposing partial, request-scoped constraints without exponential solver call blow-up.
@@
    Audit-guided fine-tuning and distribution-aware decoding. Although PRISM is model-agnostic and achieved the reported results without domain-specific fine-tuning, the framework exposes systematic opportunities to leverage audit trails for targeted model improvement. The runtime audit logs $\tau$ and AGR repair traces capture precise failure modes and correction patterns that reflect domain-specific constraint violations. These audit-derived signals can guide efficient fine-tuning in vertical domains by identifying high-value constraint patterns where model alignment would reduce repair overhead. Pre-training or fine-tuning on (UMM, ICM, $\Pi$)-annotated corpora could bias $p_{\text{LLM}}$ toward emitting structures that already satisfy $\mathcal{C}_{\mathrm{struct}}$ and respect high-value semantic invariants, reducing AGR workload and improving first-pass success rates, particularly for the $\sim$20% of architecturally dense cases that currently require iterative repair.
    A critical consideration is that grammar-constrained decoding (GCD) may induce distribution distortion by aggressively masking tokens to enforce structural constraints. Recent grammar-aligned decoding (GAD) (Park et al., 2025a) methods aim to respect grammar constraints while preserving the base model’s token distribution. Future work should explore hybrid approaches that combine audit-guided fine-tuning with distribution-aware sampling techniques. Selectively applying stricter GCD enforcement only to safety-critical structural regions while employing GAD-style preservation in descriptive regions could balance formal correctness guarantees with linguistic naturalness. The audit trail $\tau$ provides direct empirical evidence of where enforcement overhead is justified versus where relaxed sampling suffices, enabling data-driven calibration of this enforcement/naturalness trade-off.
@@
    Domain applicability and scenario recommendations. The UMM–ICM–CVG architecture exhibits varying degrees of suitability across different LLM application scenarios. For transformations from natural language or multimodal inputs to structured artifacts—the focus of this work—PRISM provides formal correctness guarantees and audit trails that address the trustworthiness requirements of regulated domains. The framework is particularly well-suited to domains characterized by explicit meta-models (or meta-models that can be inductively extracted from specifications), comprehensive constraint specifications spanning structural, semantic, and logical dimensions, and validation requirements that demand machine-checkable evidence of conformance.
@@
PRISM operationalizes evidence-carrying artifact generation for regulated domains through layered constraint enforcement (L1 structural automata, L2 SHACL/SMT validation) and first-class provenance. This UMM–ICM–CVG pattern offers a practical path toward trustworthy AI-assisted engineering and legal reasoning workflows requiring formal verification and auditability.

### v2

@@ 4.5. RQ3: Cross-Domain Transferability via UMM S2 Path
We evaluate the framework in a domain without an explicit meta-model: cross-border civil and commercial jurisdiction under the Brussels I bis Regulation. This setting probes whether the UMM S2 path can induce a usable meta-model and constraint graph directly from statutes and cases, and whether the ATLAS architecture validated in AUTOSAR can be transferred into a high-stakes legal reasoning setting that demands doctrinal correctness and faithful citation rather than domain-specific engineering artifacts.
@@ 4.5.1. Experimental Setup
We examine a legal domain without a formal schema (Brussels I bis). Prior work notes domain hallucination risks (Magesh2025); we therefore use the UMM S2 path to induce entities/relations/constraints from text and test whether layered enforcement reduces doctrinal errors (Magesh2025).
@@
Figure 10. UMM S2 pipeline architecture for cross-domain Private International Law evaluation. The pipeline instantiates a UMM mini-instance from fact patterns, links induced factors to Brussels I bis provisions through ICM construction, applies promote-aware constrained decoding with JSON Schema enforcement, and performs multi-layer verification with deterministic repair. Output comprises Evidence Registry and Structured Decision JSON forming a complete Decision Package.
Architecture of the UMM S2 pipeline for evaluating cross-domain Private International Law scenarios. Fact patterns are used to instantiate a UMM mini-instance. Induced factors are linked to Brussels I bis provisions via Integrated Constraint Model (ICM) construction. The system then applies promote-aware constrained decoding with JSON Schema enforcement, followed by multi-layer verification and deterministic repair. The final outputs are an Evidence Registry and a Structured Decision JSON, which together form a complete Decision Package.
@@
  • RAG: Retrieval-augmented prompting in the style of commercial legal assistants studied in (Magesh2025). The model receives the fact pattern plus top-$k$ Brussels I bis provisions retrieved for that scenario, but generation remains unconstrained: it may assert a jurisdictional basis even if statutory preconditions are not actually satisfied.
  • ATLAS: Full UMM $\rightarrow$ ICM $\rightarrow$ CVG S2 pipeline. The system first induces the structured mini-instance (UMM), links factors to provisions (ICM), then performs promote-aware JSON Schema-constrained decoding followed by deterministic verification and repair. This configuration enforces precedence, exclusivity, pairing requirements, and formality gates before finalizing the decision.
@@
The evaluation uses 60 expert-authored Brussels I bis jurisdiction cases spanning the full hierarchy of bases of jurisdiction: exclusive (Article 24), agreement-based (Articles 25 and 31), lis pendens / first-seised stay (Article 31), special jurisdiction (Article 7), general jurisdiction (Article 4), and appearance (Article 26). Each case was drafted and doctrinally validated by a subject-matter expert in European civil procedure. As noted above, this dataset is intended to probe architectural transfer, not to claim full coverage of real-world PIL litigation.
@@ 4.5.2. Results
Table 7 summarizes results across the three configurations. ATLAS improves correctness, citation faithfulness, structural compliance, and priority alignment relative to both Baseline and RAG. RAG improves over the Baseline in several dimensions but still fails to guarantee constraint compliance.
Table 7. Cross-domain evaluation on 60 Brussels I bis jurisdiction cases comparing Baseline (LLM-only), RAG retrieval augmentation, and the full ATLAS pipeline.
Metric Baseline RAG ATLAS
@@
ATLAS achieves perfect Schema-Pass (1.000) and Rule-OK (1.000), meaning every final output is syntactically valid and satisfies precedence, exclusivity, and pairing constraints. Promotion Accuracy (alignment between the claimed forum type and the strongest legally available basis) increases from 0.367 (Baseline) to 0.567 (RAG) and 0.933 (ATLAS), indicating that ATLAS’s promote-aware decoding and deterministic repair suppress “everything applies” answers and force adherence to the Brussels I bis priority ladder.
Abstention Quality—the ability to defer judgment when required information is missing—is 0.500 for Baseline, 0.800 for RAG, and 0.800 for ATLAS. RAG benefits here because retrieved provisions explicitly enumerate preconditions; the model can recognize that those preconditions are not satisfied and refuse to commit. ATLAS preserves that epistemic caution while additionally enforcing that abstention is required when, for example, Article 25 formality prerequisites are underspecified.
@@
ATLAS incurs higher cost due to multi-stage generation, schema-constrained decoding, and post-hoc verification. This overhead is the direct cost of guaranteeing structured legality.
@@ 4.5.3. Analysis and Discussion
These results show that the UMM–ICM–CVG architecture transfers to a non-engineering, text-governed domain (PIL jurisdiction) via the S2 inductive pathway. Without any pre-existing metamodel, the system can induce a structured mini-instance, align facts to candidate provisions, enforce hierarchical precedence, and output auditable decisions. ATLAS delivers higher Legal-Correct@1 than both Baseline and RAG (0.467 vs. 0.200 / 0.300), far higher Promotion Accuracy, and perfect Schema-Pass / Rule-OK. This indicates that layered constraint enforcement meaningfully outperforms unconstrained prompting and improves upon retrieval augmentation alone (Magesh2025).
@@ Discussion.
RAG improves correctness (0.300 vs. 0.200 Baseline) and abstention (0.800 vs. 0.500), echoing (Magesh2025), but does not enforce doctrinal validity. ATLAS’s UMM $\rightarrow$ ICM $\rightarrow$ CVG pipeline adds precedence rules, exclusivity constraints, and formality gates, yielding 0.933 Promotion Accuracy, 0.778 Citation Precision, and perfect structural legality (Schema-Pass = Rule-OK = 1.000). Each case logs UMM instance, factor-provision mappings, decoding trace, and repair actions, providing end-to-end provenance for practitioner review.
This 60-case Brussels I bis evaluation is a feasibility probe testing architectural transfer, not claiming exhaustive legal coverage or deployment readiness. Scaling requires broader doctrinal coverage, multi-jurisdictional support, and human oversight in professional workflows.
@@ 5. Conclusion and Future Work
This paper presented ATLAS, a high-assurance generation framework integrating LLMs with model-driven engineering and formal methods to produce verifiable, auditable artifacts in safety- and regulation-critical domains. ATLAS combines a Unified Meta-Model (UMM) for typed domain semantics, an Integrated Constraint Model (ICM) for constraint aggregation, and Constraint-guided Verifiable Generation (CVG) with Audit-Guided Repair (AGR) to produce verifiable artifacts $\mathfrak{A}=(a,\Pi,\varphi)$. Constraints are stratified across two layers: L1 enforces structural rules during decoding (Theorem 10), while L2 validates semantic/logic constraints post-generation with machine-checkable certificates.
@@
  • System-level semantics at scale. In RQ2, architectural drift and dangling cross-file references appear in more complex systems. ATLAS can surface these violations and drive AGR, but does not yet guarantee global semantic invariants for arbitrarily dense dependency graphs.
  • From formal to programmatic scale. While SHACL and SMT provide rigorous proofs for safety-critical properties, scaling them to thousands of routine business rules introduces significant maintenance overhead and performance bottlenecks. Future industrial implementations of ATLAS may adopt a hybrid validation strategy: employing SMT only for critical safety invariants (e.g., deadlock freedom, timing bounds) while utilizing LLMs to synthesize highly optimized programmatic validators (e.g., Java/C++ assertions) from the ICM. By applying symbol-constrained decoding over the Path S1 APIs, this approach will combine the flexibility of LLM code generation with the type-safety of MDE, resolving the scalability challenges of pure formal verification.
  • Human review remains essential. ATLAS narrows the surface that humans must review and provides $\Pi$ and $\tau$ to justify remaining issues, but it does not claim unsupervised certification. Certified deployment still requires domain experts and regulatory sign-off.
@@
We outline several directions for advancing ATLAS and the broader UMM–ICM–CVG paradigm:
@@
    Incremental semantic and logic validation. Beyond the post-artifact L2 validation described in Limitations, we aim to integrate streaming SHACL engines (yeSatLMSatisfiabilityAidedLanguage2023) and incremental SMT solvers accepting partial formulas, enabling Layer-2 feedback during generation. A key challenge is exposing partial, request-scoped constraints without exponential solver call blow-up.
@@
    Audit-guided fine-tuning and distribution-aware decoding. Although ATLAS is model-agnostic and achieved the reported results without domain-specific fine-tuning, the framework exposes systematic opportunities to leverage audit trails for targeted model improvement. The runtime audit logs $\tau$ and AGR repair traces capture precise failure modes and correction patterns that reflect domain-specific constraint violations. These audit-derived signals can guide efficient fine-tuning in vertical domains by identifying high-value constraint patterns where model alignment would reduce repair overhead. Pre-training or fine-tuning on (UMM, ICM, $\Pi$)-annotated corpora could bias $p_{\text{LLM}}$ toward emitting structures that already satisfy $\mathcal{C}_{\mathrm{struct}}$ and respect high-value semantic invariants, reducing AGR workload and improving first-pass success rates, particularly for the $\sim$20% of architecturally dense cases that currently require iterative repair.
    A critical consideration is that grammar-constrained decoding (GCD) may induce distribution distortion by aggressively masking tokens to enforce structural constraints. Recent grammar-aligned decoding (GAD) (parkGrammaralignedDecoding2025) methods aim to respect grammar constraints while preserving the base model’s token distribution. Future work should explore hybrid approaches that combine audit-guided fine-tuning with distribution-aware sampling techniques. Selectively applying stricter GCD enforcement only to safety-critical structural regions while employing GAD-style preservation in descriptive regions could balance formal correctness guarantees with linguistic naturalness. The audit trail $\tau$ provides direct empirical evidence of where enforcement overhead is justified versus where relaxed sampling suffices, enabling data-driven calibration of this enforcement/naturalness trade-off.
@@
    Domain applicability and scenario recommendations. The UMM–ICM–CVG architecture exhibits varying degrees of suitability across different LLM application scenarios. For transformations from natural language or multimodal inputs to structured artifacts—the focus of this work—ATLAS provides formal correctness guarantees and audit trails that address the trustworthiness requirements of regulated domains. The framework is particularly well-suited to domains characterized by explicit meta-models (or meta-models that can be inductively extracted from specifications), comprehensive constraint specifications spanning structural, semantic, and logical dimensions, and validation requirements that demand machine-checkable evidence of conformance.
@@
  5. 5)
    Hardware-aware embedded code generation. Beyond software architecture (AUTOSAR), ATLAS is intrinsically suited for low-level embedded firmware generation. The UMM can be instantiated from hardware description standards (e.g., IP-XACT, CMSIS-SVD) to model chip peripherals and registers. The ICM can capture hardware invariants often ignored by general LLMs, such as read/write access policies, atomic bit-field manipulations, and strict timing constraints. Applying CVG in this domain could produce ”correct-by-construction” device drivers that are mathematically guaranteed to respect the hardware specification, significantly reducing the risk of bricking devices or causing physical safety hazards.
@@
ATLAS operationalizes evidence-carrying artifact generation for regulated domains through layered constraint enforcement. Our evaluation on industrial AUTOSAR systems empirically validates the 80/20 graduated automation principle: By extending the deterministic model transformation principles of S2D2 (DualStageFramework) into the realm of generative AI, ATLAS reliably offloads the bulk of routine structural construction to LLMs, while elevating human experts from manual authors to audit-enabled reviewers for the critical “last mile” of semantic refinement. By coupling the flexibility of LLMs with the accountability of MDE, ATLAS offers a practical path toward trustworthy AI-assisted engineering where complete automation is neither feasible nor desirable.